How do you solve for moles by using the molarity and volume of an acid. The only information I have is 25mL HCl. Assume 1g/mL and the molar mass of HCl is 36.48g
1
Expert's answer
2015-11-05T09:33:41-0500
V (HCl) = 25 mL (volume)
ρ (HCl) = 1 g/ml (density)
M (HCl) = 36.5 g/mol (you can`t write 36.48, it`s wrong)
n (HCl) - ? (moles)
Solution:
First way:
ρ = m / V (mass / volume),
so m = ρ * V
m (HCl) = 1 g/ml * 25 mL = 25 g
n = m / M
n (HCl) = 25 g / 36.5 g/mol = 0.685 mol
If you need to find molarity:
с (molarity) = n / V (moles of solute / liters of solution)
molarity (HCl) = 0.685 mol / 0.025 L = 27.4 mol/L (27.4 M)

Second way (the way, which you need):
molarity (HCl) = 1 000 g/L / (36.5 g/mol) = 27.4 mol/L (27.4 M),
so n (HCl) = c (molarity) * V = 27.4 mol/L * 0.025 mol = 0.685 mol